Mrs. Johnson bought ten cans of vegetables - 2 cans of corn, 1 of beans, 5 of tomatoes, and 2 of beets. If she reaches into her cupboard and pulls out one without looking, what is the chance of it being beans?
Answer by venugopalramana(3286)

Mrs. Johnson bought ten cans of vegetables - 2 cans of corn, 1 of beans, 5 of tomatoes, and 2 of beets. If she reaches into her cupboard and pulls out one without looking, what is the chance of it being beans?
TOTAL NUMBER OF CANS = 2+1+5+2=10
NUMBER OF BEAN CANS = 1
PROBABILITY OF DRAWING BEAN CAN = 1/10